# Optimizing Large Language Models for OpenAPI Code Completion

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) and their utilization in code generation tasks have significantly reshaped the field of software development. Despite the remarkable efficacy of code completion solutions in mainstream programming languages, their performance lags when applied to less ubiquitous formats such as OpenAPI definitions. This study evaluates the OpenAPI completion performance of GitHub Copilot, a prevalent commercial code completion tool, and proposes a set of task-specific optimizations leveraging Meta’s open-source model Code Llama. A semantics-aware OpenAPI completion benchmark proposed in this research is used to perform a series of experiments through which the impact of various prompt-engineering and fine-tuning techniques on the Code Llama model’s performance is analyzed. The fine-tuned Code Llama model reaches a peak correctness improvement of 55.2% over GitHub Copilot despite utilizing 25 times fewer parameters than the commercial solution’s underlying Codex model. Additionally, this research proposes an enhancement to a widely used code infilling training technique, addressing the issue of underperformance when the model is prompted with context sizes smaller than those used during training.
@@ II State of the Art in Code Generation
As technology advances, the landscape of code generation continues to evolve, with transformer-based models and integrated solutions revolutionizing the field. This section delves into the state-of-the-art solutions in code generation, explores integrated solutions that streamline the development process, assesses methodologies for evaluating the quality of generated code, and highlights the challenges in OpenAPI autocompletion as a specific case of code generation.
@@ II-A Transformer Models for Code Generation
Transformer-based models are the most widely used type of models for Natural Language Processing (NLP) and code generation in particular [5]. The original transformer model was introduced in the famous ”Attention Is All You Need” article in 2017 [6]. The model was based on the trainable neural attention mechanism proposed by Bahdanau Dzmitry et al. [7] earlier in 2014. This mechanism presented a solution to several problems of Recurrent Neural Network (RNN) based models that were one of the primary methods for NLP at the time [8].
@@
One of the main shortcomings of RNN-based models is the inability to capture long-term dependencies in the input due to vanishing and exploding gradient problems and limited hidden state size [9]. The attention mechanism allows the model to capture dependencies between any two tokens in the input sequence, drawing the likelihood of a long-term dependency being captured the same as the likelihood of any short-term dependency.
@@
Another major drawback of RNN-based models is the sequential nature of processing which makes it impossible to efficiently parallelize the computation. A significant part of the attention computation can be done independently for each token in the input sequence which dramatically speeds up processing on modern highly parallel hardware like Graphics Processing Units (GPUs) and Tensor Processing Units (TPUs).
@@
Transformer model architecture can consist of an encoder, decoder, or a combination of both. Each of these parts leverages the attention mechanism in different ways to achieve its objectives. The role of an encoder is to analyze the input sequence and encode it into an internal representation that captures dependencies between the input tokens. This is done using a mechanism called self-attention where the attention score is calculated between each pair of the input tokens and a new value is calculated for each token as a sum of all the values in the input sequence weighted by the attention score. The role of a decoder is to autoregressively generate the output sequence leveraging a mechanism called encoder-decoder attention. In this process, each output token is generated as an attention-weighted sum of the encoded input tokens and the previously generated output tokens.
@@
The encoder-only models like BERT from Google [10], RoBERTa [11], AlBERT [12] and DeBERTa [13] only consist of an encoder and do not produce sequences as the output. Instead, these models only produce encoded representations of the input sequence that can be used for tasks like input classification, sentiment analysis, or feature extraction. The objective used during training is often not well-aligned with the inference objective, so these models usually require fine-tuning on the specific target task.
@@
The encoder-decoder models like the original transformer suggested by Ashish Vaswani et al., T5 [14] and BART from Meta [15] incorporate both encoder and decoder in their architecture. These models are capable of sequence-to-sequence (seq-to-seq) generation and are commonly used for machine translation and summarization where the input and the expected output sequences can be clearly defined.
@@
The decoder-only models like the Generative Pre-trained Transformer (GPT) family of models from OpenAI [16], PaLM from Google [17], Chinchilla from Deepmind [18], and Llama from Meta [19, 20] only consist of a decoder. This preserves the ability to analyze the input sequence and generate a sequence as the output but also eliminates the boundary between the input and the output. The latter characteristic makes this type of models more generic and often abolishes the need for fine-tuning after the initial training as most of the tasks can be represented as a natural language seq-to-seq function [21]. Compared to the encoder-decoder architectures, which are also capable of seq-to-seq generation, decoder-only architectures further streamline the training process by removing the need to define the input and output formats for the model to operate on. These models are often used for language modeling and text generation but are also capable of input classification, sentiment analysis, and machine translation if invoked with a corresponding prompt. The capabilities of the large decoder-only models can be further expanded with prompting techniques like few-shot prompting [22].
@@
Code generation from a natural language task definition might be useful for no-code or low-code engineers but code completion, where the model generates continuation for an incomplete code input, represents everyday programmers’ work much closer. Models like Codex from OpenAI [24], CodeGen from Salesforce [25, 26], and Code Llama from Meta [4] leverage decoder-only architecture for generic autoregressive code modeling. This approach not only allows to generate code completion but also supports natural language task definition to some degree. These models can be prompted with a natural language code comment to generate an implementation. This can be explained by a weak natural language signal that might be captured during training on large corpora of code with comments and code-related discussions [25].
@@ II-B Transformer Model Performance Optimization
Prompt engineering and model fine-tuning are two primary methods for large language model performance optimization. A common workflow for improving the performance of a transformer-based model in a specific domain is to use prompt engineering methods to achieve the best possible performance with the base model and then fine-tune the model on a task-specific dataset if further performance improvement is required. This research will utilize both prompt engineering and fine-tuning to improve the performance of the Code Llama model on the OpenAPI definitions generation task.
@@
Prompt engineering is based on the idea that the model can be guided to generate more accurate output by optimizing the prompt in a series of human-insight-driven experiments. [27]. Common prompt engineering techniques for instruction-tuned models include few-shot prompting, chain-of-thought prompting [28], and generated knowledge prompting [29]. For autoregressive code generation, prompt engineering techniques include embedding additional context to the prompt in the form of comments or other language-specific constructs [25] and optimizing the prompt structure and size to maximize the model’s performance and resource efficiency.
@@
Fine-tuning is the second primary method for improving the performance of large language models. The idea behind this method is that the model that has already been trained for a similar or a more broad task can be further fine-tuned for a specific task. The fine-tuning process is similar to training, but it is performed on a pre-trained model and requires significantly less data and computational resources. Two fine-tuning approaches are commonly used: full fine-tuning and Parameter-Efficient Fine-Tuning (PEFT).
@@
Full fine-tuning is performed by continuing the training process on an already pre-trained model. During this process, all the model parameters are updated just like in the original training. This means that the hardware requirements for the full fine-tuning are also similar. This approach demonstrates the potential for achieving the best results, but it also requires the most computational resources and data. The ability to change all the model parameters during full fine-tuning can also lead to catastrophic forgetting, a training failure mode where the model loses the ability to perform the original task or stops generating adequate output altogether.
@@
Parameter-efficient fine-tuning, as an alternative to full fine-tuning, is performed by freezing most of the pre-trained model parameters and only training a relatively small number of additional parameters. This approach requires much less computational resources and tends to yield comparable results to the full-finetuning [30]. The PEFT approach is also considered to be less prone to catastrophic forgetting as the majority of the pre-trained model parameters are left unchanged during the fine-tuning process. Notable PEFT techniques include adapter tuning [31] which inserts additional adapter layers between the original model layers and only updates them during training, BitFit [32] which only trains the bias-terms of the original model, and Low-Rank Adaptation (LoRA) [33] which decomposes the original parameters into low-rank matrices and only trains them. The LoRA method claims to reduce the number of trainable parameters by 10,000 times and memory requirements by 3 times while maintaining or surpassing the performance of full fine-tuning when applied to models like GPT-3 and RoBERTa.
@@ II-C Integrated Code Generation Solutions
Integrated code generation and completion solutions are end-to-end systems that combine code generation models, prompt engineering techniques, developer experience optimizations, and IDE integrations with the aim of solving a real-world problem of improving developer productivity. These solutions are often proprietary, so the details of their implementation are not available to the public. However, some solutions are open-source and can be analyzed in detail.
The most notable commercial solution is GitHub Copilot. Developed by GitHub, OpenAI, and Microsoft, this code assistant is based on the Codex model. The solution is integrated into several IDEs like Visual Studio Code, NeoVim, and JetBrains IDEs. As discovered by a reverse engineering effort [34], Copilot’s prompt-building approach is based on several language-specific heuristics. For example, for programming languages like Python, a list of imported libraries and declared functions within the file is extracted and always included in the prompt. The code assistant underwent manual evaluation in a number of studies [35, 36, 37, 38] demonstrating decent performance in benchmarks like HumanEval [24] and solving moderate complexity tasks at platforms like LeetCode. No studies were found that would evaluate the solution performance in OpenAPI completion. In this research, GitHub Copilot is evaluated as the baseline solution for the OpenAPI definitions completion task.
Other proprietary solutions worth mentioning are Tabnine [39] which accents the security of the user’s data and the code base, AWS Code Whisperer [3] which demonstrates additional expertise in AWS services and client libraries and performs additional checks for license compliance and common security vulnerabilities in the generated code, and Sourcegraph Cody [40] which demonstrates awareness of the whole user’s code base.
@@ II-D Code Generation Evaluation
The current state-of-the-art evaluation method for code generation is the HumanEval benchmark. Originally proposed by OpenAI, this benchmark was designed to evaluate Python code generation by providing a set of problems each with a description and a set of automated test cases. Later, the method was extended to support 18 additional programming languages by the MultiPL-E benchmark [53]. These methods take language semantics into account but are only applicable to a limited set of programming languages and do not support specific code semantics like OpenAPI definitions.
@@ III-A OpenAPI Completion Evaluation Criteria and Benchmarking
The proposed benchmark closely resembles the real-world usage of the solution in terms of input modeling and output evaluation. Figure 1 outlines the steps for a single evaluation iteration. The masking step, where part of the input definition is replaced with a special mask token, is responsible for modeling the input. At the completion step, the evaluated model is used to replace the mask token with the generated OpenAPI definition. The evaluation step, where the completed definition is compared to the original, is responsible for the completion result evaluation.
@@
The masking step can be seen as splitting the original OpenAPI definition into two parts (prefix and suffix) and removing some amount of the definition from between. This way, the input modeling is defined by the position of the prefix-suffix split and the size of the masked part between them.
When editing an OpenAPI definition, the user does not always add new endpoints or parameters and can expect to get a completion suggestion even when they have not finished typing the previous word or syntactic construct. This means that the prefix and suffix split should be randomly selected and not tied to a new line or a keyword. Conversely, the end of the masked sequence should be on a new line, as it is reasonable to assume that users are more likely to edit a document at the end of a line rather than in the middle of it when describing new API operations or adding parameters to existing ones.
@@ III-A1 Evaluation dataset
The dataset for the benchmark was collected from the APIs-guru definitions directory [55] which is a weekly updated collection of public OpenAPI definitions. The directory contains more than 4 000 definitions in YAML format.
@@
Adyen Management API Adyen Payment system 20 359
Amazon VPC Lattice Amazon Web Services (AWS) Cloud computing 6 709
Visma e-conomic OpenAPI Visma Accounting system 10 162
Google Display & Video 360 API Google Video hosting 16 008
Mux API Mux Video hosting 7 128
Rubrik REST API Rubrik Data security 43 540
Sinao API Sinao Accounting system 12 215
@@
The Racing API The Racing API Gambling 9 650
Severa Public Rest API Documentation Visma Project management 36 871
The benchmark was implemented with the idea of being fully automated and reproducible. This made it possible to easily evaluate the performance of many solutions and configurations. However, the fact that commercial solutions like GitHub Copilot are only available as IDE plugins and require manual interaction with the IDE imposed a limitation on the size of the evaluation dataset. To keep the manual evaluation time reasonable while still providing a representative evaluation, the dataset size was constrained to 100 test cases in total which is comparable to the size of the HumanEval benchmark with 164 code problems.

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) and their utilization in code generation tasks have significantly reshaped the field of software development. Despite the remarkable efficacy of code completion solutions in mainstream programming languages, their performance lags when applied to less ubiquitous formats such as OpenAPI definitions. This study evaluates the OpenAPI completion performance of GitHub Copilot, a prevalent commercial code completion tool, and proposes a set of task-specific optimizations leveraging Meta’s open-source model Code Llama. A semantics-aware OpenAPI completion benchmark proposed in this research is used to perform a series of experiments through which the impact of various prompt-engineering and fine-tuning techniques on the Code Llama model’s performance is analyzed. The fine-tuned Code Llama model reaches a peak correctness improvement of 55.2% over GitHub Copilot despite utilizing 25 times fewer parameters than the commercial solution’s underlying Codex model. Additionally, this research proposes an enhancement to a widely used code infilling training technique, addressing the issue of underperformance when the model is prompted with context sizes smaller than those used during training. The dataset, the benchmark, and the model fine-tuning code are made publicly available.
@@ II State of the Art in Code Generation
As technology advances, the landscape of code generation continues to evolve, with transformer-based models and integrated solutions revolutionizing the field. This section delves into the state-of-the-art solutions for code generation, explores integrated solutions that streamline the development process, assesses methodologies for evaluating the quality of generated code, and highlights the challenges in OpenAPI autocompletion as a specific case of code generation.
@@ II-A Transformer Models for Code Generation
Transformer-based models are the most widely used type of models for Natural Language Processing (NLP) and code generation in particular [5]. The original transformer model was introduced in the famous “Attention Is All You Need” article in 2017 [6]. The model was based on the trainable neural attention mechanism proposed by Bahdanau Dzmitry et al. [7] earlier in 2014. This mechanism presented a solution to several problems of Recurrent Neural Network (RNN)-based models that were one of the primary methods for NLP at the time [8].
@@
One of the main shortcomings of RNN-based models is the difficulty of capturing long-term dependencies in the input due to vanishing and exploding gradient problems and limited hidden state size [9]. The attention mechanism allows the model to capture dependencies between any two tokens in the input sequence, making the likelihood of a long-term dependency being captured the same as the likelihood of any short-term dependency.
@@
Another major drawback of RNN-based models is the sequential nature of processing which makes it impossible to efficiently parallelize the computation. A significant part of the attention computation can be done independently for each token in the input sequence which dramatically speeds up the processing on modern highly parallel hardware like Graphics Processing Units (GPUs) and Tensor Processing Units (TPUs).
@@
Transformer model architecture can consist of an encoder, decoder, or combination of both. Each of these parts leverages the attention mechanism in different ways to achieve its objectives. The role of an encoder is to analyze the input sequence and encode it into an internal representation that captures dependencies between the input tokens. This is done using a mechanism called self-attention where the attention score is calculated between each pair of the input tokens and a new value is calculated for each token as a sum of all the values in the input sequence weighted by the attention score. The role of a decoder is to autoregressively generate the output sequence leveraging a mechanism called encoder-decoder attention. In this process, each output token is generated as an attention-weighted sum of the encoded input tokens and the previously generated output tokens.
@@
The encoder-only models like BERT from Google [10], RoBERTa [11], AlBERT [12], and DeBERTa [13] only consist of an encoder and do not produce sequences as the output. Instead, these models only produce encoded representations of the input sequence that can be used for tasks like input classification, sentiment analysis, or feature extraction. The objective used during training is often not well-aligned with the inference objective, so these models usually require fine-tuning on the specific target task.
@@
The encoder-decoder models like the original transformer suggested by Ashish Vaswani et al. [6], T5 [14], and BART from Meta [15] incorporate both encoder and decoder in their architecture. These models are capable of sequence-to-sequence (seq-to-seq) generation and are commonly used for machine translation and summarization where the input and the expected output sequences can be clearly defined.
@@
The decoder-only models like the Generative Pre-trained Transformer (GPT) family of models from OpenAI [16], PaLM from Google [17], Chinchilla from Deepmind [18], and Llama from Meta [19, 20] only consist of a decoder. This preserves the ability to analyze the input sequence and generate a sequence as the output but also eliminates the boundary between the input and the output. The latter characteristic makes this type of models more generic and often abolishes the need for fine-tuning after the initial training, as most of the tasks can be represented as a natural language seq-to-seq function [21]. Compared to the encoder-decoder architectures, which are also capable of seq-to-seq generation, decoder-only architectures further streamline the training process by removing the need to define the input and output formats for the model to operate on. These models are often used for language modeling and text generation but are also capable of input classification, sentiment analysis, and machine translation if invoked with a corresponding prompt. The capabilities of the large decoder-only models can be further expanded with prompting techniques like few-shot prompting [22].
@@
Code generation from a natural language task definition might be useful for no-code or low-code engineers but code completion, where the model generates continuation for an incomplete code input, represents everyday programmers’ work much closer. Models like Codex from OpenAI [24], CodeGen from Salesforce [25, 26], and Code Llama from Meta [4] leverage decoder-only architectures for autoregressive code modeling. The fill-in-the-middle (FIM) also referred to as infilling transformation and training objective allows decoder-only models to include the context from both sides of the generated part, removing the limitation of strict left-to-right generation while maintaining the ability to output long sequences of tokens [27]. This approach not only enables context-aware code completion but also supports natural language task definition to some degree. These models can be prompted with a natural language code comment to generate an implementation. This can be explained by a weak natural language signal that might be captured during training on large corpora of code with comments and code-related discussions [25].
@@ II-B Transformer Model Performance Optimization
Prompt engineering and model fine-tuning are two primary methods for large language model performance optimization. A common workflow for improving the performance of a transformer-based model in a specific domain is to use prompt engineering methods to achieve the best possible performance with the base model and then fine-tune the model on a task-specific dataset if further performance improvement is desirable. This research will utilize both prompt engineering and fine-tuning to improve the performance of the Code Llama model on the OpenAPI definitions generation task.
@@
Prompt engineering is based on the idea that the model can be guided to generate more accurate output by optimizing the prompt in a series of human-insight-driven experiments [28]. Common prompt engineering techniques for instruction-tuned models include few-shot prompting [22], chain-of-thought prompting [29], and generated knowledge prompting [30]. For autoregressive code generation, prompt engineering techniques include embedding additional context to the prompt in the form of comments or other language-specific constructs [25] and optimizing the prompt structure and size to maximize the model’s performance and resource efficiency.
@@
Fine-tuning is the second primary method for improving the performance of large language models. The idea behind this method is that the model that has already been trained for a similar or broader task can be further fine-tuned for a specific task. The fine-tuning process is similar to training, but it is performed on a pre-trained model and requires significantly less data and computational resources. Two fine-tuning approaches are commonly used: full fine-tuning and Parameter-Efficient Fine-Tuning (PEFT) [31].
@@
Full fine-tuning is performed by continuing the training process of an already pre-trained model. During this process, all the model parameters are updated just like during the original training. This means that the hardware requirements for the full fine-tuning are also similar. This approach demonstrates the potential for achieving the best results, but it also requires the most computational resources and data. The ability to change all the model parameters during full fine-tuning can also lead to catastrophic forgetting, a training failure mode where the model loses the ability to perform the original task or stops generating adequate outputs altogether.
@@
Parameter-efficient fine-tuning, as an alternative to full fine-tuning, is performed by freezing most of the pre-trained model parameters and only training a relatively small number of additional parameters. This approach requires much less computational resources and tends to yield comparable results to the full-finetuning [31]. The PEFT approach is also considered to be less prone to catastrophic forgetting as the majority of the pre-trained model parameters are left unchanged during the fine-tuning process. Notable PEFT techniques include adapter tuning [32] which inserts additional adapter layers between the original model layers and only updates them during training, BitFit [33] which only trains the bias-terms of the original model, and Low-Rank Adaptation (LoRA) [34] which decomposes the original parameters into low-rank matrices and only trains them. The LoRA method claims to reduce the number of trainable parameters by 10 000 times and memory requirements by 3 times while maintaining or surpassing the performance of full fine-tuning when applied to models like GPT-3 and RoBERTa.
@@ II-C Integrated Code Generation Solutions
Integrated code generation and completion solutions are end-to-end systems that combine code generation models, prompt engineering techniques, developer experience optimizations, and IDE integrations to solve a real-world problem of improving developer productivity. These solutions are often proprietary, so the details of their implementation are not available to the public. However, some solutions are open-source and can be analyzed in detail.
The most notable commercial solution is GitHub Copilot. Developed by GitHub, OpenAI, and Microsoft, this code assistant is based on the Codex [24] model. The solution is integrated into several IDEs like Visual Studio Code, NeoVim, and JetBrains IDEs. As discovered by a reverse engineering effort [35], Copilot’s prompt-building approach is based on several language-specific heuristics. For example, for programming languages like Python, a list of imported libraries and declared functions within the file is extracted and always included in the prompt. The code assistant underwent manual evaluation in several studies [36, 37, 38, 39] demonstrating decent performance in benchmarks like HumanEval [24] and solving moderate complexity tasks at platforms like LeetCode. No studies were found that would evaluate the solution performance in OpenAPI completion. In this research, GitHub Copilot is evaluated as the baseline solution for the OpenAPI definitions completion task.
Other proprietary solutions worth mentioning are Tabnine [40] which emphasizes the security of the user’s data and the code base, AWS Code Whisperer [3] which demonstrates additional expertise in AWS services and client libraries and performs additional checks for license compliance and common security vulnerabilities in the generated code, and Sourcegraph Cody [41] which demonstrates awareness of the whole user’s code base.
@@ II-D Code Generation Evaluation
One of the current state-of-the-art evaluation methods for code generation is the HumanEval benchmark [24]. Originally proposed by OpenAI, this benchmark was designed to evaluate Python code generation by providing a set of problems each with a description and a set of automated test cases. Later, the method was extended to support 18 additional programming languages by the MultiPL-E benchmark [54]. These methods take language semantics into account but are only applicable to a limited set of programming languages and do not support specific code semantics like OpenAPI definitions.
@@ III-A OpenAPI Completion Evaluation Criteria and Benchmarking
The proposed benchmark aims to closely resemble the real-world usage of the solution in terms of input modeling and output evaluation. Figure 1 outlines the steps for a single evaluation iteration. The masking step, where part of the input definition is replaced with a special mask token, is responsible for modeling the input. To equalize the evaluation conditions for all the solutions, the same set of masked test cases is used for all the evaluations. At the completion step, the evaluated model is used to replace the mask token with the generated OpenAPI definition. The evaluation step, where the completed definition is compared to the original, is responsible for the completion result evaluation.
@@
The masking step can be seen as splitting the original OpenAPI definition into two parts (prefix and suffix) and removing some amount of the definition from in between. This way, the input modeling is defined by the position of the prefix-suffix split and the size of the masked part between them. When editing an OpenAPI definition, the user does not always add new endpoints or parameters and can expect to get a completion suggestion even when they have not finished typing the previous word or syntactic construct. This means that the prefix and suffix split should be randomly selected and not tied to a newline or a keyword. Conversely, the end of the masked sequence should be on a newline, as it is reasonable to assume that users are more likely to edit a document at the end of a line rather than in the middle of it when describing new API operations or adding parameters to existing ones. In this study, the masked part is always 10 lines long (till the end of the 10th line), which is a reasonable size for a single parameter or a response definition.
@@ III-A1 Evaluation dataset
The dataset for the benchmark was collected from the APIs-guru definitions directory [56] which is a weekly-updated collection of public OpenAPI definitions. The directory contains more than 4 000 definitions in YAML format.
@@
Adyen API Adyen Payments 20 359
Google Video API Google Video 16 008
Mux API Mux Video 7 128
Rubrik API Rubrik Security 43 540
Severa Public API Visma Project mngmt. 36 871
Sinao API Sinao Accounting 12 215
@@
The Racing API The Racing API Sports 9 650
VPC Lattice API AWS Cloud 6 709
Visma e-conomic Visma Accounting 10 162
The benchmark was implemented with the idea of being fully automated and reproducible. This made it possible to easily evaluate the performance of many solutions and configurations. However, the fact that commercial solutions like GitHub Copilot are only available as IDE plugins and require manual interaction with the IDE imposed a limitation on the size of the evaluation dataset. To keep the manual evaluation time reasonable while still providing a representative evaluation, the dataset size was constrained to 100 test cases in total (10 test cases from each of the 10 definitions) which is comparable to the size of the HumanEval benchmark with 164 code problems.